# Self-Calibrated CLIP for Training-Free Open-Vocabulary Segmentation

### v2

@@ 4.3 Ablation Study
We conduct ablation study of the multi-level fusion strategy in Table 6 with following approaches. ①baseline: $\mathbf{L}(\mathbb{X}^{penul})$direct sum: $\mathbf{L}(\mathbb{X}^{penul})+\sum_{i\in\mathcal{M}}\mathbb{X}^{i}$one pass: $\mathbf{L}(\mathbb{X}^{penul}+\sum_{i\in\mathcal{M}}\mathbb{X}^{i})$two pass: $\mathbf{L}(\mathbb{X}^{penul})+\mathbf{L}\left(\sum_{i\in\mathcal{M}}\mathbb{X}^{i}\right)$. Here, $\mathbf{L}$ is the last layer, and $\mathcal{M}$ is the multi-level features. The results align with our analysis in Section 3.4, validating two key principles: (1) using the last layer for alignment: comparison between ② and ④ reveals that ④ aligns the multi-level features using the parameter space of the last layer, which improves feature compatibility and leads to a significant performance boost. (2) maintaining the integrity of the last feature: comparison between ③ and ④ indicates that ③ weakens the original last feature, resulting in a performance drop. Additionally, we compare our approach with the dual-path strategy proposed in CLIPSurgery [29], where each layer of the image encoder in the new path is modified to self-self attention, and the outputs of the new path are directly aggregated. This can be formulated as: dual path: $\mathbf{L}(\mathbb{X}^{penul})+\sum_{i\in\mathcal{M}}\mathbb{X}^{i}_{new}$, where $\mathbb{X}^{i}_{new}$ is the feature from the new path. This strategy can be regarded as an enhanced version of ②. However, in comparison, our method achieves superior results while avoiding the additional computational overhead introduced by the new path.
To analyze the impact of selected layers in multi-level fusion, we report the average mIoU across eight datasets in Table 7, indicating that our method is robust to the choice of selected layers, as different combinations yield similar results. The notation 4–10 in the table denotes fusing all features from layer 4 to 10, which achieves the best 43.9% mIoU, surpassing the baseline (w/o) by 1.5% mIoU.
Efficiency Comparison
As shown in Table 8, we present the efficiency comparison on the COCO-Object dataset using a V100 GPU. Compared to the previous state-of-the-art method, ProxyCLIP [26], SC-CLIP requires no additional backbones, significantly improving efficiency. Specifically, the FPS increase from 3.9 to 6.5, and the FLOPs are reduced from 34.4G to 17.5G. Additionally, we show the detailed breakdown of the computational cost for each component.
Table 8: Efficiency comparison of training-free methods.
@@
+MultiFuse 17.5 149.6 6.5
5 Visualization
In Figure 6, we present the qualitative results of various methods, showing that SC-CLIP consistently achieves more precise and higher quality results than other methods. In contrast, CLIP’s results are considerably noisy and exhibit a homogeneity effect (e.g., in the 2nd row, the entire image is labeled as “dog”). By performing self-calibration, SC-CLIP enables CLIP to produce semantically coherent results.
6 Conclusion
@@
We present Self-Calibrated CLIP (SC-CLIP), a training-free method to enhance the open-vocabulary segmentation performance of CLIP. We first identify and resolve the negative impact of anomaly tokens. Then we leverage the semantic consistency within CLIP’s intermediate features to enhance both feature discriminability and attention correlation. Together with multi-level fusion strategy, SC-CLIP sets new state-of-the-art results, without requiring additional data, parameters or backbone. This demonstrates the inherent properties of CLIP can be effectively leveraged to calibrate itself to produce semantically coherent representations.

### v3

@@ IV-C Ablation Study
Two-pass Strategy: We conduct the ablation study of the multi-level fusion strategy in Table IX with the following approaches. 1 baseline: $\mathbf{L}(\mathbb{X}^{penul})$ 2 direct sum: $\mathbf{L}(\mathbb{X}^{penul})+\sum_{i\in\mathcal{M}}\mathbb{X}^{i}$ 3 one pass: $\mathbf{L}(\mathbb{X}^{penul}+\sum_{i\in\mathcal{M}}\mathbb{X}^{i})$ 4 two pass: $\mathbf{L}(\mathbb{X}^{penul})+\mathbf{L}\left(\sum_{i\in\mathcal{M}}\mathbb{X}^{i}\right)$. Here, $\mathbf{L}$ is the last layer, and $\mathcal{M}$ is the multi-level features. The results align with our analysis in Section III-D, validating two key principles: (1) using the last layer for alignment: comparison between 2 and 4 reveals that 4 aligns the multi-level features using the parameter space of the last layer, which improves feature compatibility and leads to a significant performance boost. (2) maintaining the integrity of the last feature: comparison between 3 and 4 indicates that 3 weakens the original last feature, resulting in a performance drop. Additionally, we compare our approach with the dual-path strategy proposed in CLIPSurgery [39], where each layer of the image encoder in the new path is modified to self-self attention, and the outputs of the new path are directly aggregated. This can be formulated as: dual path: $\mathbf{L}(\mathbb{X}^{penul})+\sum_{i\in\mathcal{M}}\mathbb{X}^{i}_{new}$, where $\mathbb{X}^{i}_{new}$ is the feature from the new path. This strategy can be regarded as an enhanced version of 2. However, in comparison, our method achieves superior results while avoiding the additional computational overhead introduced by the new path.
TABLE XI: Efficiency comparison of training-free methods.
@@
+TwoPass 17.5 149.6 6.5
TABLE XII: Generalization to other ViT-based backbones.
Method V21 C60 Obj C59 Stf Avg
MetaCLIP [72] 56.4 32.1 31.3 36.0 24.2 36.0
+Ours 63.7 36.9 37.3 40.4 26.9 41.0
OpenCLIP [13] 55.8 31.4 31.8 34.8 23.4 35.4
+Ours 60.7 35.2 36.3 39.2 27.3 39.7
SigLIP [82] 27.1 11.5 10.9 18.4 12.0 16.0
+Ours 38.3 19.2 18.0 24.1 15.9 23.1
BLIP [36] 51.1 28.3 31.2 29.4 20.2 32.0
+Ours 58.7 34.8 36.7 37.3 24.9 38.5
To analyze the impact of selected layers in multi-level fusion, we report the average mIoU across eight datasets in Table X, indicating that our method is robust to the choice of selected layers, as different combinations yield similar results. The notation 4–10 in the table denotes fusing all features from layer 4 to 10, which achieves the best 43.9% mIoU, surpassing the baseline (w/o) by 1.5% mIoU.
Efficiency Comparison: As shown in Table XI, we compare the efficiency of different methods on the COCO-Object dataset [7] using an NVIDIA V100 GPU. Compared to the previous state-of-the-art method, ProxyCLIP [34], our SC-CLIP eliminates the need for additional backbones, leading to a significant improvement in efficiency. Specifically, the FPS increases from 3.9 to 6.5, and the FLOPs are reduced from 34.4G to 17.5G. We also provide a detailed breakdown of the computational cost for each component. By default, the LOF algorithm is implemented with NumPy, which is computationally inefficient. As discussed in Section III-B, we re-implement LOF using PyTorch, resulting in a substantial inference speed-up. In particular, the FPS improves from 6.3 to 7.6, while maintaining functional consistency.
IV-D Discussion
Figure 7: Detailed analysis of other ViT-based backbones.
Figure 8: Qualitative results of improvement analysis, where each strategy is progressively added to demonstrate its effect.
Figure 9: PCA analysis of the ResNet-based CLIP.
In this section, we provide a broader discussion on the applicability and generalization of our method. We begin by evaluating its applicability to different ViT-based vision–language models, assessing the consistency and robustness of our approach across models that vary in training data and strategies. We then turn to ResNet-based CLIP models to examine whether the anomaly token issue also arises in non-ViT architectures. Finally, we explore the potential of our approach in open-vocabulary 3D perception tasks, demonstrating its applicability to more complex scenarios. Together, these analyses provide a more comprehensive validation for our method.
Different ViT-based Vision–Language Models: To further examine the generalization of our method, we evaluate it on several ViT-based vision–language models, which all adopt image–text contrastive pretraining but differ in certain aspects. Specifically, MetaCLIP [72] and OpenCLIP (pretrained on LAION) [13] vary in the pretraining data, while BLIP [36] and SigLIP [82] introduce alternative training paradigms or architectures. For each model, we adopt the same configurations as those used for CLIP in Section IV-A. As shown in Table XII, our method consistently improves performance across all cases, showing its generalization to diverse models.
To further explain why our method generalizes well across different models, we provide analyses in Figure 7 (better viewed when zoomed in). In Figure 7 (a), using the image from Figure 3 for illustration, the attention maps of four different selected patch tokens (marked by $\bigstar$) show that different tokens all exhibit excessive focus on the same regions, corresponding to anomaly tokens highlighted by the orange dashed circles. In Figure 7 (b), PCA analysis of patch-level features reveals these anomaly tokens as clear outliers, further confirming their presence across all models. In Figure 7 (c), ROC curves comparing mid-layer and last-layer features show that mid-layer features consistently achieve higher AUC values across all models, indicating stronger discriminative ability. Building on this observation, our method explicitly addresses anomaly tokens and leverages mid-layer spatial consistency, leading to consistent relative AUC improvements, with gains of 19.1% for MetaCLIP, 18.8% for OpenCLIP, 21.2% for BLIP, and 12.5% for SigLIP. Overall, these results explain the root of our method’s generalization ability: by tackling the consistently observed anomaly token issue and enhancing local correspondence, it achieves robust improvements across diverse ViT-based vision–language models.
ResNet-based CLIP Models: We further examine ResNet-based CLIP variants (ResNet-50, ResNet-101, and ResNet-50x4) to investigate whether the anomaly token issue also exists beyond ViT architectures. As shown in the PCA visualizations in Figure 9, the feature distributions of ResNet variants are tightly clustered within a narrow range (e.g., [3,3][-3,3]). In contrast, ViT-based models (Figure 7 (b)) exhibit clear outliers, with x-axis values reaching as high as 100–1000. Quantitative results in Table XIII further confirm this observation: resolving anomaly tokens (denoted as “+AnomRes”) leads to negligible changes in average performance. These findings indicate that ResNet-based CLIP does not suffer from the anomaly token issue. We hypothesize that this is due to the strong local inductive bias of convolutional architectures, which prevents the extreme outliers that often arise in ViT-based models.
Open-vocabulary 3D perception tasks: Open-vocabulary 3D perception has recently gained increasing attention, aiming to extend vision–language models to tasks such as 3D instance segmentation and object detection, with representative approaches including OpenMask3D [60], OpenIns3D [27], Coda [8], and INHA [29]. To further validate the applicability of SC-CLIP in 3D scenarios, we investigate the open-vocabulary 3D semantic segmentation task using OpenScene [50]. In this setting, the model is required to perform per-point classification, which strongly relies on dense feature representations and naturally aligns with the design of SC-CLIP. Specifically, we employ the image feature fusion method (corresponding to the “2D fusion” setting in their paper), which enables a direct evaluation of the visual encoder without additional training. Experiments on the ScanNet [16] validation set and the Matterport3D [11] test set, as reported in Table XIV, show that SC-CLIP achieves the best performance among all training-free methods. The remaining performance gap compared with training-based methods such as LSeg [35] and OpenSeg [21] can be largely attributed to their additional supervised training of both the image encoder and decoder. Overall, these results demonstrate that SC-CLIP not only advances open-vocabulary 2D segmentation, but also generalizes effectively to 3D scenarios, thereby further validating its versatility and applicability.
Figure 10: Qualitative Results of Open-Vocabulary Segmentation. We compare our method with CLIP [52], MaskCLIP [86], SCLIP [62] and ClearCLIP [33], all without post-processing. Our SC-CLIP produces much clearer and more accurate results.
Figure 11: Failure case analysis in textual and visual aspects.
TABLE XIII: Resolving anomalies in ResNet-based CLIP.
Method V20 C60 Obj C59 Stf Avg
ResNet50 57.5 6.4 6.6 9.8 5.7 17.2
+ AnomRes 57.5 6.5 6.6 9.8 5.7 17.2
ResNet101 61.8 7.5 7.2 10.3 5.8 18.5
+ AnomRes 61.7 7.6 7.4 9.8 5.4 18.4
ResNet50x4 57.6 6.7 6.3 9.1 4.9 16.9
+ AnomRes 57.6 7.1 6.8 8.8 4.9 17.0
TABLE XIV: Performance comparison of different methods on open-vocabulary 3D semantic segmentation.
Methods ScanNet [16] Matterport3D [11]
mIoU mAcc mIoU mAcc
training-based methods
LSeg [35] 50.0 62.7 32.3 40.0
OpenSeg [21] 41.4 63.6 32.4 45.0
training-free methods
CLIP [52] 1.5 5.1 1.5 4.2
SCLIP [62] 15.8 40.3 14.8 28.5
ClearCLIP [33] 21.5 52.6 22.0 40.8
ProxyCLIP [34] 22.6 53.4 23.7 42.0
SC-CLIP (Ours) 25.1 55.5 25.1 42.7
IV-E Visualization
In Figure 8, we present a qualitative improvement analysis which clearly illustrates the progressive improvements introduced by each strategy and highlights the overall contributions of our method. Compared to the baseline result, resolving the anomaly tokens effectively removes the noise in the lower part of the image (d vs. c). Further applying the self-adjusting strategy enhances local correspondence and leads to better overall coherence, as can be observed in the upper region of the result (e vs. d). Finally, incorporating the two-pass strategy provides richer details, such as a more complete structure of the bicycle frame and the seat (f vs. e).
Beyond the improvement analysis, we further provide a qualitative comparison across different representative methods in Figure 10, including CLIP [52], MaskCLIP [86], SCLIP [62], ClearCLIP [33], and our SC-CLIP. The results demonstrate that SC-CLIP (last column) consistently produces more precise and higher-quality outcomes compared to the other methods, showing robustness across different visual contexts. In contrast, CLIP’s results are considerably noisy and suffer from a homogeneity effect (e.g., in the second row, the entire image is labeled as “dog”). Previous approaches such as SCLIP and ClearCLIP focused only on modifying the attention computation, without explicitly resolving anomaly tokens or enhancing local correspondence, which makes their results still suboptimal. By directly addressing these issues, our method achieves more semantically consistent outputs with sharper boundaries and finer structural details.
IV-F Failure Case Analysis
We further analyze the failure cases of our method, as illustrated in Figure 11. Our motivation is to enhance CLIP’s local correspondence in a training-free manner; however, the approach is inherently bounded by CLIP’s representational capacity. Specifically, the limitations manifest in two aspects: (1) Textual aspect. As shown in Figure 11 (a), when using generic input prompts such as “Tree” or “Person”, the segmentation results are suboptimal, producing vague or incomplete masks. In contrast, more specific and descriptive inputs like “Tree With Snow” or “Skiing Person” lead to significantly improved results. This phenomenon suggests that for CLIP-based models, using more specific and descriptive inputs leads to better results; (2) Visual aspect. The limited resolution and training paradigm of the CLIP image encoder make it difficult to accurately represent small objects or boundary regions in complex scenes. As shown in Figure 11 (b), such limitations result in coarse boundaries (e.g., the contour of the car on the left) and recognition errors (e.g., failing to detect the distant traffic sign). These failure cases provide a more comprehensive perspective on the applicability and limitations of our method.
V Conclusion
@@
In this paper, we present Self-Calibrated CLIP (SC-CLIP), a training-free approach designed to enhance the open-vocabulary segmentation performance of CLIP. We observe that anomaly tokens induce uniform attention patterns and feature homogenization, which compromise spatial representation. To address this, we mitigate their adverse effects from two complementary perspectives. First, we explicitly identify anomaly tokens and replace them based on local context. Second, we reduce their influence on normal tokens by enhancing feature discriminability and attention correlation, leveraging the intrinsic semantic consistency embedded in CLIP’s mid-level features. Together with a two-pass strategy, SC-CLIP achieves state-of-the-art performance without requiring additional data, parameters, or backbones. These results demonstrate that the inherent capabilities of CLIP can be effectively leveraged to calibrate itself and produce semantically coherent representations.
The motivation of this work is to enhance the dense feature representation of CLIP while preserving its original capabilities in a training-free manner. For future work, we believe that to fundamentally improve the model, modifications to the architecture and training strategy are necessary. Naturally, this would necessitate retraining the model from scratch, along with substantial computational resources and large-scale data.
Acknowledgment
This work was supported in part by the Shenzhen Science and Technology Program under Grant CJGJZD20220517142402006, and National Natural Science Foundation of China (Grant No. 62206153).